# A Survey on Text-guided 3D Visual Grounding: Elements, Recent Advances, and Future Directions

### v1

@@ 2.4 Scene and Text Context Encoder
Figure 4: Illustration of two types of scene context encoders. It learns more dependencies among the intra-modal instances (objects or points).
The common scene context encoders can be categorized into two main types: the approaches [50, 47] based on graph neural networks [51] and the approaches [23, 27] based on self-attention mechanisms [52], as illustrated in Fig. 4. Scene encoders based on graph neural networks are mostly employed in the two-stage T-3DVG framework. It explicitly establishes a graph structure over the object proposals obtained through 3D detectors. Each node on the graph represents an object proposal and each edge represents the relationship between the proposals. Typically, this graph is constructed based on the spatial and semantic distances between proposals, with edges connecting nearby proposals. After constructing the proposal graph, Graph Convolutional Networks [51] or other similar graph neural networks [53, 54] are employed to encode the relationships between objects into the features of the proposals. Besides, scene encoders based on self-attention mechanisms represent another popular approach that can handle both object-level and point-level visual features. Unlike graph neural network-based encoders, self-attention mechanisms operate without an explicit graph structure, allowing for not only automatic semantic correlation learning among object proposals but also complex spatial and relational capturing for the whole point cloud. There are some scene encoders [25] based on self-attention mechanisms also incorporate attention between object features and point features to further refine object features from the points contained within them. Some T-3DVG methods [55, 56] also combine the advantages of these two designs, which manually inject spatial relationships between objects into self-attention mechanisms through methods such as position encoding or attention modulation. Additionally, some methods [30, 57, 58] propose to enrich proposal information by adding multi-view information or encoding ternary relationships. Overall, the general scene context encoding process can be described as:
@@
Text Context Encoder. Common text context encoders can be categorized into two types according to existing T-3DVG methods: the approaches [8, 7, 50] based on recurrent neural networks and the approaches [23, 27, 28] utilizing self-attention mechanisms. As for the first type of approach, their popular recurrent neural network encoders include GRU [59] and LSTM [60], which effectively capture temporal dependencies in text. These recurrent neural network encoders can further be classified into two categories based on their textual output: word-level encoders and sentence-level encoders. Word-level encoders maintain the dimensionality corresponding to the number of words in their output, while sentence-level encoders encode an entire sentence into a single vector feature as output. The following illustrates both kinds of encoders:
@@
As for the second type of approach, they utilize general self-attention [52] that is capable of capturing long-range dependencies in sentences more efficiently, albeit potentially requiring greater computational resources. There are some methods [27] that directly utilize a combination of these two types of encoders for contextual learning. Other methods also employ more powerful large-model encoders such as CLIP [61] for semantic learning. As described in Section 2.3, some methods also utilize pre-trained language models like RoBERTa [46] as text feature extractors, which also includes encoding of textual context relationships. Overall, the encoder using self-attention or pre-trained model is typically at the word level, and its computational process can be represented as follows:
@@
2.5 Multi-Modal Interaction
@@
Some early T-3DVG methods [8, 7, 62, 50, 47] directly employ a simple concatenation operation to coarsely combine scene features and text features that have already been embedded in the same space. Considering such modality fusion based on concatenation cannot finely integrate and match the detailed features of the two modalities, some post-processing is usually required to merge the features of both modalities into the same feature space. Common post-processing methods include MLP [7] and GNN [8, 50, 47], which are utilized to further refine the fused representations. There are also some methods that utilize attention modules or large models to achieve this. This types of multi-modal interaction module can be represented as follows:
@@
In addition to the multi-modal attentive fusion, these methods are also coupled with individually designed pre-processing modules. The main purpose of this pre-processing is to generate better attentive queries, attentive keys, and attentive values in the attention mechanism. A common pre-processing approach involves an initial filtering of objects or points, selecting those closely related to the query text based on simple criteria. This filtering can be achieved through explicit filtering based on object categories [65] or through implicitly neural network-based learning [25, 64]. Some other methods [66] employ multi-modal alignment mechanisms as pre-processing, which aligns information from the two modalities at different levels to enable the cross-attention mechanism to better fuse language and visual information. This types of multi-modal interaction module can be represented as follows:
@@
2.6 Grounding Head and training objectives
@@
2.6.1 Proposal Ranking Technique
For methods using a two-stage framework, since the location of all possible objects has been obtained through a detector or segmentor, it is only necessary to sort the matching degree between these objects and the text through a Proposal Ranking Technique. Given the object location $\mathcal{D}\in\mathbb{R}^{M\times 6}$ obtained by the detector or segmentor, these methods use a classifier to calculate a confidence score $\mathbf{s}\in\mathbb{R}^{M}$ for each object based on the fused features $\mathbf{C}_{\text{fus}}$, representing the matching degree between the object and the query text, that is, the possibility that a certain object is the target object. The classifier is usually implemented using an MLP with a sigmoid function. Finally, the location of the object $\mathcal{D}_{j}\in\mathbb{R}^{6}$ with the highest confidence score $j=\operatorname*{arg\,max}(\mathbf{s})$ is output as the grounding result.
2.6.2 Proposal Regression Technique
@@ 2.6.3 Training objectives
As for the proposal-based technique, the primary training objective is the object grounding loss $\mathcal{L}_{\text{VG}}$, which is usually implemented using the cross-entropy function to supervise the confidence score for each possible object:
@@
Figure 7: Depiction of a typical instance from scene $0$_00 within the ScanRefer Dataset [7]. This dataset is recognized for its thorough object labeling methodology, where almost every object in the scene is annotated and provided with five associated descriptions.
In addition to the primary training objective of object grounding, many methods [8, 7] also design some auxiliary tasks for robust representation learning, such as object detection, text classification, etc.. The object detection loss is determined by the detector or segmentor used. For example, for methods that use VoteNet [33] as the detector, the object detection loss $\mathcal{L}_{\text{det}}$ is obtained by weighted summation of the vote regression loss $\mathcal{L}_{\text{vote}}$ [33], objectness binary classification loss $\mathcal{L}_{\text{obj}}$, box regression loss $\mathcal{L}_{\text{box}}$, and the semantic classification loss $\mathcal{L}_{\text{sem-cls}}$:
@@
As for the proposal-free technique, similar losses in the proposal-based technique are also employed to supervise feature learning. The main difference is that the proposal-free technique only requires the confidence scoring loss function and a regression loss $\mathcal{L}_{\text{regress}}$ [25] to supervise the quality of the generated object bbox.
@@
3 Datasets and Measures
3.1 Benchmark Datasets

### v2

@@ II-D Scene and Text Context Encoder
Figure 4: Illustration of two types of scene context encoders. They learn more dependencies among intra-modal instances (objects or points).
The common scene context encoders can be categorized into two main types: the approaches [56, 53] based on graph neural networks [57] and the approaches [29, 33] based on self-attention mechanisms [58], as illustrated in Fig. 4. Scene encoders based on graph neural networks are mostly employed in the two-stage T-3DVG framework. It explicitly establishes a graph structure over the object proposals obtained through 3D detectors. Each node on the graph represents an object proposal and each edge represents the relationship between the proposals. Typically, this graph is constructed based on the spatial and semantic distances between proposals, with edges connecting nearby proposals. After constructing the proposal graph, Graph Convolutional Networks [57] or other similar graph neural networks [59, 60] are employed to encode the relationships between objects into the features of the proposals. Besides, scene encoders based on self-attention mechanisms represent another popular approach that can handle both object-level and point-level visual features. Unlike graph-neural-network-based encoders, self-attention mechanisms operate without an explicit graph structure, allowing for automatic semantic correlation learning among object proposals. Some scene encoders [31] based on self-attention mechanisms also incorporate attention between object features and point features to further refine object features from the points contained within them. Some T-3DVG methods [61, 62] also combine the advantages of these two designs, which manually inject spatial relationships between objects into self-attention mechanisms through methods such as position encoding or attention modulation. Additionally, some methods [36, 63, 64] propose to enrich proposal information by adding multi-view information or encoding ternary relationships. Overall, the general scene context encoding process can be described as:
@@
Text context encoder. Common text context encoders can be categorized into two types according to existing T-3DVG methods: the approaches [12, 11, 56] based on recurrent neural networks and the approaches [29, 33, 34] utilizing self-attention mechanisms. As for the first type of approaches, their popular recurrent neural network encoders include GRU [65] and LSTM [66], which effectively capture temporal dependencies in the text. These recurrent neural network encoders can further be classified into two categories based on their textual output: word-level encoders and sentence-level encoders. Word-level encoders maintain the dimensionality corresponding to the number of words in their output, while sentence-level encoders encode an entire sentence into a single vector feature as output. The following illustrates both kinds of encoders:
@@
As for the second type of approaches, they utilize general self-attention [58] that is capable of capturing long-range dependencies in sentences more efficiently, albeit potentially requiring greater computational resources. There are some methods [33] that directly utilize a combination of these two types of encoders for contextual learning. Other methods also employ more powerful large-model encoders such as CLIP [67] for semantic learning. As described in Section II-C, some methods also utilize pre-trained language models like RoBERTa [52] as text feature extractors, which also includes encoding of textual context relationships. Overall, the encoder using self-attention or pre-trained model is typically at the word level, and its computational process can be represented as follows:
@@
II-E Multi-Modal Interaction
@@
Some early T-3DVG methods [12, 11, 68, 56, 53] directly employ a simple concatenation operation to coarsely combine scene features and text features that have already been embedded in the same space. Considering such modality fusion based on concatenation cannot finely integrate and match the detailed features of the two modalities, some post-processing is usually required to merge the features of both modalities into the same feature space. Common post-processing methods include multilayer perceptron (MLP) [11] and graph neural networks (GNN) [12, 56, 53], which are utilized to further refine the fused representations. There are also some methods that utilize attention modules or large models to achieve this. This type of multi-modal interaction module can be represented as follows:
@@
In addition to the multi-modal attentive fusion, the existing methods are also coupled with individually designed pre-processing modules. The main purpose of this pre-processing is to generate better attentive queries, attentive keys, and attentive values in the attention mechanism. A common pre-processing approach involves an initial filtering of objects or points, selecting those closely related to the query text based on simple criteria. This filtering can be achieved through explicit filtering based on object categories [71] or through implicitly neural network-based learning [31, 70]. Some other methods [72] employ multi-modal alignment mechanisms as pre-processing, which aligns information from the two modalities at different levels to enable the cross-attention mechanism to better fuse language and visual information. This types of multi-modal interaction module can be represented as follows:
@@
II-F Grounding Head and Training Objectives
@@
II-F1 Proposal Ranking Technique
For methods using a two-stage framework, since the location of all possible objects has been obtained through a detector or segmentor, it is only necessary to sort the matching degree between these objects and the text through a Proposal Ranking Technique. Given the object location $\mathcal{D}\in\mathbb{R}^{M\times 6}$ obtained by the detector or segmentor, the existing methods use a classifier to calculate a confidence score $\mathbf{s}\in\mathbb{R}^{M}$ for each object based on the fused features $\mathbf{C}_{\text{fus}}$, representing the matching degree between the object and the query text, that is, the possibility that a certain object is the target object. The classifier is usually implemented using an MLP with a sigmoid function. Finally, the location of the object $\mathcal{D}_{j}\in\mathbb{R}^{6}$ with the highest confidence score $j=\operatorname*{arg\,max}(\mathbf{s})$ is output as the grounding result.
II-F2 Proposal Regression Technique
@@ II-F3 Training Objectives
As for the two-stage technique, the primary training objective is the object grounding loss $\mathcal{L}_{\text{VG}}$, which is usually implemented using the cross-entropy function to supervise the confidence score for each possible object:
@@
In addition to the primary training objective of object grounding, many methods [12, 11] also design some auxiliary tasks for robust representation learning, such as object detection, text classification, etc.. The object detection loss is determined by the detector or segmentor used. For example, for methods that use VoteNet [39] as the detector, the object detection loss $\mathcal{L}_{\text{det}}$ is obtained by weighted summation of the vote regression loss $\mathcal{L}_{\text{vote}}$ [39], objectness binary classification loss $\mathcal{L}_{\text{obj}}$, box regression loss $\mathcal{L}_{\text{box}}$, and the semantic classification loss $\mathcal{L}_{\text{sem-cls}}$, with balanced weights $\lambda$ as:
@@
As for the one-stage technique, similar losses in the two-stage technique are also employed to supervise feature learning. The main difference is that the one-stage technique only requires the confidence scoring loss and a regression loss $\mathcal{L}_{\text{regress}}$ [31] to supervise the quality of the generated object.
Figure 7: The taxonomy of T-3DVG based on different methods. The fully-supervised methods and weakly-supervised methods will be introduced in Sections III.A and III.B. The methods of using other modality input and large models will be described in Sections III.C and III.D.
@@
III Methods
Most of the existing T-3DVG methods are proposed with the fully-supervised learning paradigm. Early approaches primarily utilize pre-trained object detectors or segmentation models to pre-sample all potential objects from the complicated 3D scene. Then, these object proposals are then paired with a given sentence query to learn their semantic correspondence through cross-modal interaction to determine the most matching one. However, this two-stage framework not only severely relies on the quality of the 3D detectors and segmentation models, but also requires time-consuming reasoning among the multiple proposals. To alleviate this limitation, without using object proposals, the one-stage framework is proposed to take the whole 3D scene as the input and embed point-level features in an end-to-end manner. After point-text cross-modal reasoning, it directly regresses the spatial bounding box of the text-guided object. Although the above two types of fully-supervised methods have achieved significant performance in recent years, they require a large number of bounding boxes for numerous objects in the 3D point cloud scenes, along with their descriptive texts, to provide reliable supervision. Considering manually annotating these object-text pairs is very time-consuming and labor-intensive, some works propose to solve the T-3DVG task in the weakly-supervised setting, which only utilizes text labels for grounding without relying on any object bounding box annotations. There are also some recent works that utilize other-modal input (instead of point cloud) for grounding or incorporate traditional grounding models with powerful large language models.
In this section, we provide a complete review of existing T-3DVG methods via a taxonomy in Fig. 7, based on the methods’ architectures and learning algorithms, to better categorize T-3DVG approaches. Specifically, we discuss the unique characteristics of each category and illustrate the highlights of each T-3DVG method in details.
III-A Fully-Supervised Methods
III-A1 Two-Stage Methods
Most existing T-3DVG methods [11, 12, 71, 56, 33, 34, 53, 27, 28, 29, 61, 78, 62, 79, 72, 80, 81, 64, 82, 69, 83, 84, 85, 86, 35, 63, 87, 88, 89, 50, 90, 91, 92] are deployed in a two-stage framework, which first utilizes a pre-trained 3D detection or segmentation model to obtain all possible 3D objects in the scene and then individually matches and reasons them with the query for selecting the best one. These works can be further categorized into a basic proposal-then-selection approach and a multi-modal enhanced proposal-then-selection approach. In particular, the latter approach integrates more contextual modalities’ information into the 3D reasoning process for enriching the representation learning.
Basic proposal-then-selection methods. ScanRefer [11] (as shown in Fig. 8) and ReferIt3D [12] are the pioneer methods proposed to address the T-3DVG task following a basic proposal-then-selection methods. They generally first utilize a pre-trained 3D object detector to obtain all possible objects within the whole 3D scene. Then, they match the objects’ features with the semantic of text descriptions for determining the best aligned object as the final output. Instead of using pre-trained 3D objectors, a few works [71, 56] also introduce to utilize a pre-trained segmentor to obtain fine-grained semantic features for text matching. Considering the object detector is relatively lightweight and enough to capture object-level features, almost all two-stage methods follow the detection-based framework.
Figure 8: ScanRefer architecture, reproduced from Chen et al. [11].
Besides, there are also some works [27, 28, 29, 79, 64, 82, 85, 94, 95] designed to address the task-specific challenges in T-3DVG. Considering that object locations are important to reason the spatial relations mentioned in the sentence, spatial embeddings [27, 28] are designed to embed positional-encoded bounding box information into the attention mechanism. To help distinguish complicated background objects from the foreground target, He et al. [29] propose to separately align low-level and high-level of words with the objects for comprehending and recognizing the corresponding object entity and relation. Kato et al. [79] introduce the ARKitSceneRefer method to tackle the object location on small indoor 3D objects. It introduces a new small-object dataset and develops a simple coarse-to-fine framework to achieve the goal. Hsu et al. [64] propose to translate language into programs with hierarchical structures by leveraging large language-to-code models and extend prior neuro-symbolic visual reasoning methods by introducing functional modules that effectively reason about high-arity relations, key in disambiguating objects in complex 3D scenes. This framework can alleviate two issues: the expense of labeling and the complexity of 3D grounded language. Zhang et al. [82] introduce a new case when multiple objects exist in the same scene, and devise a contrastive learning strategy to distinguish the objects for better determining. Considering unstructured natural utterances and scattered objects might lead to undesirable performances, Wu et al. [85] propose to leverage language models to parse language description, suggesting a referential order of anchor objects. Such ordered anchor objects allow the model to update visual features and locate the target object during the grounding process. Instead of localizing a unique pre-existing object within a single 3D scene like previous works, Huang et al. [95] introduce a more realistic setting, named Group-wise 3D Object Grounding, to simultaneously process a group of related 3D scenes, allowing a flexible number of target objects to exist in each scene.
Moreover, some proposal-then-selection works [61, 78, 62, 72, 69, 83] design novel training strategies or relevant joint tasks to assist 3D grounding. For example, Cai et al. [61] and Chen et al. [83] propose to jointly train a 3D captioner and a 3D grounder to provide diverse query descriptions for better supervising the grounding module with mutual information learning. Based on this joint framework, Chen et al. [78] further incorporate a teacher-student architecture into the model design to train a robust grounder by mimicking the strong ability of the joint framework. Liu et al. [94] propose to utilize few-shot annotations (e.g., 10% data) to train the grounding models with full supervision. Specifically, they develop a teacher-student framework with an additional captioning module to generate and correct pseudo labels for training. Zhu et al. [62] propose a pre-training strategy with transformers for 3D vision and text alignment, which utilizes self-attention layers for both single-modal modeling and multi-modal fusion without any sophisticated task-specific design. Specifically, it collects diverse prompts, scene graphs, 3D scans, and objects with self-supervised pre-training. Jin et al. [72] also propose a 3D vision-language pertaining strategy with context-aware spatial-semantic alignment and mutual 3D-language masked modeling to learn more representative features. It shows great performance on the 3D grounding task. Unal et al. [69] develop three novel stand-alone modules to improve grounding performance for challenging repetitive instances, i.e., instances with distractors of the same semantic class. Two new datasets [96, 97] are proposed to scale the knowledge learning-ability of T-3DVG.
Multi-modal enhanced proposal-then-selection methods. Although the basic proposal-then-selection methods achieve significant performance, they solely exploit the contexts within the original 3D point cloud input for retrieving the specific objects according to different queries, lacking sufficient yet complementary information from other views or modalities for enriching the retrieval knowledge. To this end, there are many recent methods [35, 63, 87, 88, 98, 99, 36, 50, 89, 90, 91] proposed to further perceive more contextual information from other perspectives (such as 2D images, multi-view contexts, etc.) for enriching the traditional 3D grounding knowledge, as shown in Fig. 9.
Considering the traditional point cloud data are sparse, noisy, and contain limited semantic information compared with 2D images, many works [35, 63, 87, 88, 89, 100, 90] propose to utilize additional 2D image input to assist 3D visual grounding. These methods generally first detect potential object proposals through 3D object detector, then encode each object’s 2D image into feature vectors to combine with 3D object features for aligning with the semantics of sentence description. Since more information is utilized, these works achieve better performance than basic proposal-then-selection frameworks. In addition to the point cloud input, Xiao et al. [98] propose to build a new representation for 3D points in the semantic point cloud defined on high-level semantics. This new representation guides the model to understand more relationships by simplified representation of objects, rather than limited to traditional appearance learning. Both semantic and appearance object features are concatenated to feed into the latter cross-modal reasoning module. To make T-3DVG more practical in real-world applications, Lu et al. [99] introduce a human-robot interaction fashion that takes human guidance as a cue to facilitate the development of 3D visual grounding. A point cloud of a human agent is specially introduced as additional input along with a designed gesture extraction module. Therefore, the object-level scene representations are reasoned with both human guidance and textual semantics to determine the location of target object. Besides, multi-view learning designs [36, 50, 91] is also a promising multi-modal enhancement strategy. They take multi-view point clouds of the same 3D scene as inputs, and mine their correspondences between the same object for grasping and comprehending the view knowledge. Overall, these multi-modal enhanced two-stage methods achieve better performance as more contextual information is embedded into the representation learning.
III-A2 One-Stage Methods
Without relying on the quality of pre-trained object generators (i.e., 3D detector or segmentor), some recent T-3DVG methods follow the one-stage framework that trains the 3D grounding models end-to-end from the feature extraction to the final cross-modal grounding. Among them, inspired by the already successful two-stage approach, some works [30, 101, 102, 103, 104, 105, 106, 76, 77] propose a unified detection-interaction method that jointly trains both object detection and object reasoning with mutual learning. To reduce the training parameters and resources, recent works [31, 32, 73, 107, 75, 108] directly encode the point-wise features of the whole point cloud and globally match all points with the query to regress the optimal locations of object boundary box locations.
Figure 10: Uni3DL architecture, reproduced from Li et al. [102].
Figure 11: 3D-SPS architecture, reproduced from Luo et al. [31].
Based on the powerful grounding ability of 3D-SPS and EDA, the latter works [73, 107, 109] focus more on the detailed network designs. Wang et al. and Jain et al. [73, 75] emphasize the importance of relative spatial relationships among objects described by the sentence description, and propose a 3D relative position module to learn and reason the spatial relations between the key point features before regressing their corresponding bbox outputs. Huang et al. [30] develop a contextual global transformer with learnable queries to iteratively generate text-described objects in the 3D scene. Then, it develops a fine-grained local transformer to associate the object features for scoring and selection. This paper designs more efficient regression module within the grounding pipeline, achieving end-to-end training and reducing the resource costs. Moreover, unlike previous pre-set objects containing lots of background objects, the generated objects of this paper are text-related, significantly improving the bbox quality. Lin et al. [107] incorporate the superpoint mask segmentation into the object grounding task with a mutual learning framework for joint training, bringing additional mask-level knowledge for improving the grounding performance. Wang et al. [108] design contextual geometric feature modeling and complex utterance understanding with further semantic-geometric consistency learning, providing sufficient knowledge for regressing the object bbox. Compared to unified detection-interaction methods, these regression-based methods are more efficient.
Figure 12: WS-3DVG architecture, reproduced from Wang et al. [37].
III-B Weakly-Supervised Methods
Wang et al. [37] propose the first weakly-supervised framework for T-3DVG task, as shown in Fig. 12. To fill the information gap between 3D visual features and textual features, it first extracts all possible objects from the whole 3D scene with a pre-trained 3D detector. Then, it coarsely selects the top-$K$ object proposals based on their semantic similarities with the sentence description. To further mine their potential relationships, it borrows the text-based pre-training strategies from transformer [58], which masks crucial words in the sentence and reconstructs the masked keywords using each selected proposal one by one. The reconstructed accuracy finely reflects the semantic similarity of each object to the query. Therefore, the best matched object is chosen as the final grounding result. Xu et al. [110] also introduce a 3D-VLA network to address the weakly-supervised T-3DVG task. It leverages additional 2D semantics to assist the grounding process. Specifically, it encodes object-level features with pre-trained 3D detector and image-text features with CLIP [67]. Then, it interacts 3D-level, image-level, and text-level semantics to determine which object is matched with the textual description. In summary, there are a few works that make attempts to investigate the challenging weakly-supervised T-3DVG problem. However, their frameworks are still limited without special designs of robust representation learning, resulting in worse performance than fully-supervised methods.
III-C Methods with Other-Modality Input
Instead of using a point cloud as the 3D grounding input, there are also a few works [68, 111, 112] proposed to address the task from a new perspective with other types of inputs, such as RGB-D images and human-interaction vectors. Liu et al. [68] propose to ground objects in single-view RGB-D images. Given an RGB-D image and a sentence query input, they first fuse the multi-modal features via attention mechanisms and generate a visual heatmap to coarsely localize the relevant regions in the RGB-D image. Then, an adaptive feature learning based on the heatmap is introduced to perform the object-level matching to finely locate the referred object. Different from previous point-cloud-based grounding, this RGBD-based grounding requires fewer resources to collect and annotate data. Another RGBD-based method [111] is also proposed to handle the holistic multi-modal 3D perception suite. In addition to the RGB-D input, this method also feeds corresponding RGB images into the model. The whole network encodes both 2D and 3D visual features to densely interact with sentence queries, leading to more fine-grained object grounding. Instead of relying on RGB-D data input, Zhan et al. [112] propose to ground target objects in monocular images. As we know, the task of monocular object detection [113, 114] task is well developed and has a large amount of collected data, providing more robust training. Besides, PATRON [104] takes multiple views of non-verbal interactions to locate the target objects in more practical robotics scenarios. Lin et al. [115] also introduce a wild case of automatic driving data and corresponding 2D images for grounding. Overall, the above methods provide a new perspective for T-3DVG instead of understanding and reasoning point clouds.
Figure 13: LAMM architecture, reproduced from Yin et al. [101].
III-D Methods using Large Models
By benefiting from the strong comprehension of large language models (LLMs), there are many recent T-3DVG methods [116, 112, 50, 106, 117, 80, 85, 101, 118, 97, 119, 120, 121] that adopt large models to construct their grounding framework. Considering the description annotations are complicated but with limited information, most works [116, 112, 50, 106] utilize LLM tools [122, 123, 124, 125, 119] to decompose complex natural language queries into semantic constituents for adaptive grounding, and enrich the contexts of original description with designed prompts for providing more detailed commonsense knowledge, or directly generate a new description from another perspective via captioners. Yang et al. [116] further utilize LLMs to evaluate the spatial relations among the proposed objects to make a final grounding decision. These methods incorporated more diverse textual semantics into the text-object reasoning process, leading to more precise grounding results. Besides, some works [117, 80, 85] attempt to leverage the commonsense reasoning capability of LLMs to make inferences about the objects in the scene and their interactions. Specifically, Bakr et al. [80] and Wu et al. [85] propose to exploit Chain-of-Thoughts [126, 127] to decompose the referring task into multiple interpretable steps, whereas to reach the final target the model must first predict the anchors one by one, in a logical order. However, the above methods simply take the LLMs as a module tool to assist the traditional grounding model, failing to exploit the strong power of the LLM architecture designs. To this end, recent methods [101, 118, 97, 121] extend the reasoning ability of multi-modal LLMs [128, 129, 130] into the 3D scenarios, as shown in Fig. 13. Specifically, they design object or scene-level 3D encoders to embed the visual input into the same latent vectors as the language dimensions in LLMs. Then, both textual and visual embeddings are jointly fed into the LLMs for reasoning. The grounding results can be achieved by inputting a specific prompt into the LLMs.
III-E Discussion
In summary, existing T-3DVG methods can be generally categorized into two types: fully-supervised approaches and weakly-supervised approaches. As for fully-supervised approaches, most works follow a basic proposal-then-selection framework, which first utilizes pre-trained models to extract all possible object proposals in each scene and then match them with textual sentences for selection. To enhance the prior knowledge for assisting grounding, some proposal-then-selection methods also utilize multi-modal data as additional input to provide sufficient grounding information. Although this two-stage framework achieves significant performance, it severely relies on the quality of the pre-trained model and costs large resources to compute background objects. Therefore, one-stage methods are proposed to directly regress the target objects. Some of them follow the two-stage methods to design a unified detection-interaction frameworks, which replace the pre-trained object model with an efficient learnable one. The others directly regress the target object with point-level feature reasoning. As for weakly-supervised approaches, they solely exploit the knowledge of the relationship annotations between scenes and queries without object-level bounding boxes. Existing weakly-supervised works generally follow a proposal-then-matching framework that scores potential proposals with self-learning strategies. There are also some works that utilize other 3D-modal input instead of point clouds, or combine LLMs with traditional grounding models, or propose to address outdoor scenes [131, 132]. Overall, there is still plenty of room for exploration in developing methods of weakly-supervised approaches and methods of using large models.
IV Datasets and Measures
IV-A Benchmark Datasets
Figure 14: Depiction of a typical instance from scene_id $0$_00 within the ScanRefer Dataset [11]. This dataset is recognized for its thorough object labeling methodology, where almost every object in the scene is annotated and provided with five associated descriptions.
Achieving optimal performance in the T-3DVG task involving 3D vision and language heavily relies on the availability of extensive datasets. However, the process of acquiring and annotating volumetric data presents formidable hurdles, resulting in smaller collections of RGB-D datasets [133, 134, 135, 136] for 3D scenarios compared to their 2D counterparts. Consequently, researchers often resort to employing synthetic data to supplement the dearth of authentic data from the real world [137, 138]. To tackle this challenge, Dai et al. pioneered the creation of the ScanNet dataset [38], an extensively annotated compilation of 3D indoor scans derived from real-world settings, captured through a proprietary RGB-D capture mechanism.
Comprising 2.5 million perspectives derived from 1513 scans, drawn from 707 unique environments, the ScanNet dataset surpasses other well-known datasets [139, 140, 134, 141, 136] like NYUv2 [139], SUN3D [134], and SceneNN [136] in terms of scale. Nevertheless, it is essential to recognize that ScanNet primarily caters to tasks such as categorizing 3D objects, labeling semantic voxels, and retrieving 3D objects with annotations at the instance level within a 3D point cloud. Consequently, it may not be directly applicable to tasks like 3D visual grounding and 3D dense captioning. To address this constraint, novel datasets such as ScanRefer [11], Nr3D [12], and Sr3D [12] have emerged, specifically tailored for 3D visual grounding tasks, building upon the foundation laid by ScanNet. Both ScanRefer and Nr3D/Sr3D curate subsets of scenes from ScanNet, accompanied by additional manually crafted descriptions for the objects within each scene. Moreover, they provide accessible online platforms for data visualization, rendering them invaluable assets for the development and assessment of models in the realm of 3D visual grounding tasks.